A tea of 45°C was emptied by Tunde. If the temperature at half cup level was 30°C after 2 minutes. How long did he take to empty the cup if the final temperature was 20°C. Take the temperature of the surrounding as 15°C.
1
Expert's answer
2020-05-03T17:23:24-0400
"T(x)=A+Be^{-\frac{x}{X}}, A=15"

"T(0)=45=A+B=15+B\to B=30"

"T(2)=15+30e^{-\frac{2}{X}}=30\to 30e^{-\frac{2}{X}}=15"

"X=2.8854\ min"

Thus,


"20=15+30e^{-\frac{x}{2.8854}}"

"x=5.17\ min"